OpenStudy (anonymous):

is 5x^-2y^-1 a polynomial?

OpenStudy (compassionate):

\[5x^? \] What is 5x to the power of?

OpenStudy (anonymous):

5x^-2

OpenStudy (anonymous):

-2

OpenStudy (jhannybean):

\[\huge (5x^{-2y})^{-1}\]

OpenStudy (anonymous):

there isnt any parentheses

OpenStudy (jhannybean):

What is this, multivariable calculus? The graph is represented by a 3-dimensional space...

OpenStudy (jhannybean):

There isn't parenthesis, but to properly show that a function is raised twice parenthesis are needed, in LaTeX format.

OpenStudy (anonymous):

it is a question asking if it is a polynomial...I don't think it is because there is no degree...I am dumb

OpenStudy (anonymous):

Thanks :)

OpenStudy (jhannybean):

Haha no problem.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

wait can you help me with one more? :)

OpenStudy (anonymous):

How do I divide x^8+x^4/x?

OpenStudy (anonymous):

is 2 questions not legal?

OpenStudy (jhannybean):

Lol as long as someone is willing to help you,sure :P but if theyre wanting you to post up another question so everyone could have an equal chance to answer, then you should do that ^_^ \[\large \frac{x^8+x^4}{x}\] divide each part of the numerator by "x" \[\large \frac{x^8}{x}+\frac{x^4}{x}\] are you able to solve the rest? :)